OpenStudy (anonymous):

The function f(x) = x². The graph of g(x) is f(x) translated to the right 3 units and down 3 units. What is the function rule for g(x)? (confused more than words can say)

jimthompson5910 (jim_thompson5910):

Hint: if you translate f(x) using those rules to get g(x), you will get g(x) = f(x-3) - 3

jimthompson5910 (jim_thompson5910):

Example: f(x) = x^2 + 2x + 5 f(x-3) = (x-3)^2 + 2(x-3) + 5 ... translate 3 units to the right f(x-3) - 3 = (x-3)^2 + 2(x-3) + 5 - 3 ... translate 3 units down f(x-3) - 3 = (x-3)^2 + 2(x-3) + 2 So g(x) = (x-3)^2 + 2(x-3) + 2